OpenStudy (anonymous):

How to integrate (x^2 -1)/x? Do I use u-substitution?

OpenStudy (anonymous):

what can you do here? Can you use algebra to make this problem easier before you integrate?

OpenStudy (anonymous):

do you mean long division?

OpenStudy (anonymous):

nah just divide each term by x

OpenStudy (anonymous):

\[\frac{x^2-1}{x}=x-\frac{1}{x}\]

OpenStudy (anonymous):

try splitting up the fraction...as sat as done

OpenStudy (anonymous):

oh I see. so in the end the answer would be \[1/2x ^{2}-\ln \left| x \right|+ c?\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

good, but always try and look for algebra tricks before doing the integral It will help you in the long run!

OpenStudy (anonymous):

yes, after you told me told me that I felt really dumb. i'll definitely try to do that in the future.

OpenStudy (anonymous):

good you will do fine. also just because someone just easily gives you the answer doesn't mean you shouldn't try and figure it out for yourself.

OpenStudy (anonymous):

thanks for the advice. sometimes i want to quit but i really want do well on the bc exam. so i keep trying to understand what im doing. you are so encouraging, seriously.